OpenStudy (oswaldmurphy):

A polynomial function can be written as (x − 2)(x − 3)(x + 5). What are the x-intercepts of the graph of this function? (2, 0), (3, 0), (−5, 0) (−2, 0), (−3, 0), (5, 0) (2, 0), (3, 0), (5, 0) (−2, 0), (−3, 0), (−5, 0) Would this be A?

OpenStudy (oswaldmurphy):

@welshfella

OpenStudy (oswaldmurphy):

@.Sam. @MARC_D @ganeshie8

OpenStudy (welshfella):

the x intercepts occur when f(x) = 0 so we can write (x - 2)(x - 3)(x + 5) = 0 what are the values of x which satisfy the above?

OpenStudy (welshfella):

Note that if any of these factors are equal to 0 than the whole function equals zero.

OpenStudy (welshfella):

So what value of x makes x - 2= 0?

OpenStudy (oswaldmurphy):

-2

OpenStudy (welshfella):

no for x - 2 to be = 0 x must be 2 (2-2=0)

OpenStudy (welshfella):

now you need to find what value of x in (x - 3) and also (x + 5) make them = 0.

OpenStudy (oswaldmurphy):

Yea sorry I am stupid XD

OpenStudy (oswaldmurphy):

Ok, so it is A?

OpenStudy (oswaldmurphy):

Thats what I thought before, just checking and kinda wanted an explanation. Thanks!

OpenStudy (welshfella):

x + 3 = 0 so x = 3 x + 5 = 0 so x = -5

OpenStudy (welshfella):

as -5 + 5 = 0

OpenStudy (welshfella):

yes A is correct

OpenStudy (oswaldmurphy):

Thank you!

OpenStudy (welshfella):

yw